# Rethinking Diffusion for Text-Driven Human Motion Generation

### v1

@@ 2.3 Impact on Method Evaluation Robustness
The widely adopted evaluators [11] utilize all features including redundant 3 components which can be imprecise and unfair.
Table 3: Existing Evaluator performance results with altering HumanML3D data by adding noise or replacing with noise in essential and redundant dimensions. The evaluator heavily emphasizes redundant dimensions during evaluation process.
@@
Table 4: The evaluation results using evaluators trained on all vs. essential dimensions on HumanML3D data. VQ-based models perform similarly with both evaluators, while diffusion-based models are disadvantaged when evaluated with redundancy.
@@
In Tab. 4, we compare model performance when evaluated on all versus essential dimensions. Results indicate that VQ-based methods exhibit more consistent error across dimensions, benefit from all-dimension evaluation. Diffusion-based methods, despite excelling in essential dimensions, are penalized for fluctuation in redundant dimensions. This underscores a more robust evaluator is crucial.
@@ 3.1 Reforming Motion Representations
To systematically address the limitations of motion representations in Sec. 2.2, we use only the essential feature groups (i.e., the first #$\#\text{joints}\times 3+1$ dimensions). After excluding the redundant dimensions, we avoid mixing representations from various distributions, such as 6D rotational and categorical. The retained features are all 3D continuous representations, ensuring a uniform distribution that aligns better with diffusion-based generation framework.
To further optimize the motion representations, we then project those essential features into a compact and fine-grained latent space using a motion AutoEncoder (AE). Compared with the motion Variational AutoEncoder (VAE) [5], the deterministic AE projection avoids the variation in the motion latents (i.e. $\boldsymbol{\epsilon}$ incorporation), providing more stable representations that are better suited for diffusion modeling and motion reconstruction.
@@
Following previous works using latent space [13, 43, 5], the encoder $E$ downsamples $\mathbf{X}$ from $N$ length to $\mathbf{x}$ of $n$ length. The decoder $D$ upsamples it back to $N$ length. With this integration, our method can also use motion latent $\mathbf{x}$ in the diffusion process, which offers acceleration for both training and sampling. In addition, since the downsampling introduces temporal awareness [11], it can improve the temporal coherence of autoregressive diffusion in Sec. 3.2.
@@
where $t$ is the continuous timestep. Then the velocity for diffusion to predict is defined as:
@@
where $\alpha_{t}$, $\sigma_{t}$ are continous time coefficients, $\dot{\alpha}_{t}=\frac{\mathbf{d}\alpha_{t}}{\mathbf{d}t}$, $\dot{\sigma}_{t}=\frac{\mathbf{d}\sigma_{t}}{\mathbf{d}t}$. We can also deduce the score $\mathbf{s}(\mathbf{x},t)$ from $\mathbf{v}(\mathbf{x},t)$ if needed.
@@ 3.2 Diffusion: An Autoregressive Approach
Autoregressive methods [13, 30, 31, 43] demonstrate significant advantages in motion generation. Instead of generating the entire sequence $\mathbf{x}^{1:n}$ with a condition c by modeling $p(\mathbf{x}^{1:n}|c)$, autoregressive generation defined as “predicting next token/tokens with previous token/tokens”, simplify the generation process as sequentially generating continuations through chains:
@@
where each conditional probability $p(\mathbf{x}^{i}|\mathbf{x}^{<i})$ represents the likelihood of generating motion $\mathbf{x}_{i}$ given previous ones $\mathbf{x}^{<i}$.
Naively training diffusion models to perform autoregressive generation using MSE loss between the model output and ground truth fails as it simplifies to a regression problem rather than explicitly capturing the chained probabilistic distributions of $p(\mathbf{x}^{1:n}|c)$ in autoregressive generation.
Inspired by recent image generation advances [22, 40], which demonstrate the potential of autoregressive continuous image modeling by leveraging logits from an autoregressive model as conditioning parameters into a continuous sampling network to better model underlying probability, we revisit human motion diffusion models from an autoregressive perspective, leveraging its advantages of simplified generation process as chained continuations.
@@ 3.2.1 Masked Autoregressive Motion Generation
We follow the masked autoregressive approach proposed by MAR [22]. For human motion generation, in each autoregressive iteration, we define unmasked motion latents as $\text{um}=\mathbf{x}^{i_{1}:i_{k}}$ and masked motion latents as $\text{m}=\mathbf{x}^{j_{1}:j_{n-k}}$. The unmasked latents can be refined in a new pseudo-order $\text{um}=\mathbf{x}^{\prime 1:k}$ to serve as previously generated blocks, and the masked motion latents $\text{m}=\mathbf{x}^{\prime k+1:n}$ represent the motion latents that need to be generated based on um with condition $c$. Formally, this process is $p(\mathbf{x}^{\prime 1:n}|c)$ =
$p(\text{m}|c)\prod_{j=1}^{k}p(\text{m}|\text{um})\vspace{-1.5mm}$ (13)
where $k$ is number of autoregressive steps needed to produce the entire sequence. we also visually illustrate motion masked autoregressive in the right-most part of Fig. 4.
@@ 3.2.2 Autoregressive Generation Branch
Masked Autoregressive Transformer is designed to process time-variant motion data and provide rich contextual condition $z$ for the diffusion branch. Given unmasked motion latent sequence um as previous generated motion latents, the masked autoregressive transformer $g$ will produce conditions $z$ to the diffusion branch to generate latents at position of masked motion latents m by:
@@
Given the relative simplicity of motion data compared to image data, we choose to incorporate a single AdaLN [14] transformer [42] layer and leverage the generation performance and speed benefits from bidirectional attention shown in many previous methods [31, 30, 13].
Diffusion MLPs adopts MLPs as its primary structure because with autoregressive modeling, the motion data input into the diffusion branch is independent single $D$-dimensional motion latents. The single D-dimensional structure input data aligns well with the MLP’s simplicity and strength in channel-wise manipulation. Unlike MAR [22], where scaling focuses on understanding models, in motions, we scale diffusion MLPs to various model sizes. Using the condition $z$ from Masked Autoregressive Transformer, diffusion branch produce each motion latents $\mathbf{x}^{\prime i}$ in masked token m’s position at each timestep $t$ by:
@@
During training, we randomly mask a subset of $k$ motion latents with a learnable continuous mask vector following the cosine masking schedule from MoMask [13]. The autoregressive model learns to provide accurate signals on the unmasked latents given time step and text condition, with the diffusion MLPs learn to utilize this signal to predict $\boldsymbol{\epsilon}$ or $\mathbf{v}(\mathbf{x},t)$. Training objectives for entire generation branch denote as:
@@
modeling noise prediction. For velocity prediction, $\mathcal{L}_{GB}=$
@@
During sampling, given previous latents um, we simply add mask vectors to the sequence, allowing the autoregressive model to generate appropriate signals for masked positions and feed into the diffusion MLPs to sample latents. The sampling process can be denoted as:
@@
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ for noise prediction. For velocity prediction using ODE sampling with step size $\Delta t$ then:
@@
To address biases in the current evaluation approach shown in Sec. 2.3, we propose a new evaluation framework that focuses exclusively on essential features, enabling a fairer generation evaluation across different-based methods.
We first construct an evaluator that retains the architecture of the widely used evaluator [11], consisting of a convolutional movement encoder, a GRU[6]-based motion encoder, and a GRU-based text encoder using GloVe[29] embeddings, to ensure consistency. This evaluator is trained solely on essential dimensions to prioritize features that meaningfully contribute to final motion generation.
@@
By training exclusively on essential dimensions, we ensure the evaluators capture only meaningful features from the final generated motion. The adoption of dual evaluators also provides a more robust and comprehensive framework to accurately compare different-based generation methods.
@@ 4 Experiment
Evaluation metrics Following Section 3.3, we employ two evaluators trained with only essential dimensions: one architecturally identical to the one proposed in T2M [11] and a CLIP-based evaluator. Using the T2M evaluator, we adopt evaluation metrics from T2M, including (1) R-Precision (Top-1, Top-2, and Top-3 accuracies) and Matching, which measures the semantic alignment between generated motion embeddings and their corresponding captions’ glove embedding; (2) Fréchet Inception Distance (FID), which assesses the statistical similarity between ground truth and generated motion distributions; and (3) MultiModality, which measures the diversity of generated motion embeddings per same text prompt. Using the CLIP-based evaluator, we include the CLIP-score [16], which measures the compatibility of motion-caption pairs by calculating the cosine similarity between generated motion embedding and motion caption.
@@ 4.2 Results and Analysis
Following previous works [11, 39], we conduct each experiment 20 times on both datasets and report the mean result along with a 95% confidence interval. To ensure a fair comparison on the new evaluators, we train all baseline models from scratch following their original methods on the same obtained dataset (full dimension). We present the quantitative results of our method alongside baseline state-of-the-art human motion generation methods in Tab. 5, and the qualitative comparison results in Fig. 5. In addition we also present model scalability results in App. C
As observed, our method achieves superior performance across multiple metrics, including FID, R-Precision, Matching score, and CLIP-score, consistently outperforming baseline methods with non-marginal improvements on both KIT-ML and HumanML3D datasets. Compared to diffusion-based baseline methods, our approach showcases a significantly stronger ability to generate stable motion that follows closely to text instructions and aligns with ground truth. Notably, while the SOTA diffusion-based baseline method ReMoDiffuse [45] relies on additional data retrieval from a large database to achieve high-quality motion generation, our method surpasses ReMoDiffuse’s performance results without requiring auxiliary formulations. In comparison to VQ-based baseline methods, our approach maintains better motion quality and also delivers greater diversity, a quality where VQ-based methods typically underperform.
@@ 4.3 Ablation Study
In the ablation study, we further study the impact of reforming data representation and distribution, and autoregressive modeling in our method. We present the ablation result in Tab. 6. The results demonstrate that both these proposed components contribute greatly.
@@ 5 Related Work
VQ-Based Human Motion Generation TM2T [12] first introduces Vector Quantization (VQ) to text-to-human motion generation, enabling discrete motion token modeling. T2M-GPT [43] extended this by leveraging a GPT[3] to motion autoregressive generation. Subsequent methods have sought to integrate a larger model[19, 46] (e.g. large language models), or manipulate attention mechanisms  [48]. Most recently, MMM [31] and MoMask [13] revisit generation methodology by employing bidirectional attention-based masked generation techniques inspired by MaskGIT [4]. BAMM [30] introduced a dual-iteration framework that combines unidirectional generation with bidirectional refinement to enhance the coherence of generated motions. In this paper, we examine the strengths of these approaches and improve a diffusion model inspired by these insights.
Diffusion-Based Human Motion Generation. Inspired by the success of denoising diffusion models in image generation domain [18, 37], several pioneering works [39, 20, 44] have adapted denoising diffusion processes to human motion generation. Building on these works, MLD [5]further optimized the denoising process in latent space to improve training and sampling efficiency. Recent methods have diversified their focus, exploring retrieval-augmention [45], controllable generation [7], as well as investigating advanced architectures [47] such as Mamba [9]. In this paper, we thoroughly investigate the limitations of diffusion-based methods and propose a novel approach to address them.
Autoregressive Generation with Continous Data. GIVT first introduced the idea of leveraging outputs from an autoregressive model as parameters for a Gaussian Mixture Model, enabling probabilistic autoregressive modeling and generation. MAR then utilized logits from a masked autoregressive model as input to a small diffusion branch, producing more fine-grained generation. Inspired by these approaches, in this paper, we propose a novel framework that integrates diffusion-based motion generation with autoregression, combining the strengths from both worlds to achieve enhanced generative performance.
@@ 6 Conclusion
In conclusion, we introduce a novel diffusion-based generative framework for text-driven 3D human motion generation. Our method reforms motion data representation and distribution to better fit the diffusion model, incorporated bidirectional masked autoregressive training and sampling techiniques, and is evaluated by more robust evaluators. Extensive experiments demonstrate our method’s superior motion generation performance across evaluation metrics in both KIT-ML and HumanML3D datasets.

### v2

@@ 2.3 Impact on Method Evaluation Robustness
The widely adopted evaluators [24] utilize all features including redundancy which is imprecise and unfair.
Table 3: The result with existing evaluator on HumanML3D dataset. We alter data by adding noise or replacing it with noise in essential and redundant dimensions. The result shows the evaluator heavily emphasizes redundant dimensions during evaluation.
@@
Table 4: The evaluation results using evaluators trained on all vs. essential dimensions on HumanML3D. VQ-based models significantly outperform diffusion-based models under all-dimension evaluation, but gap closes under essential evaluation.
@@
In Tab. 4, we compare model performance when evaluated on all versus essential motion dimensions. VQ-based methods (e.g., T2M-GPT) tend to outperform diffusion-based ones (e.g., MDM) under all-dimension evaluation, benefiting from more consistent error across dimensions. However, when focusing only on animation-relevant essential dimensions, MDM performs comparably to T2M-GPT, aligning better with visual perception. This suggests that diffusion-based methods, despite excelling in essential dimensions, may be penalized by fluctuations in redundancy and underscores a more robust evaluator is crucial.
@@ 3.1 Reforming Motion Representations
To systematically address the limitations of motion representations in Sec. 2.2, we only use the essential feature groups (i.e., the first #$\#\text{joints}\times 3+1$ dimensions). After excluding the redundant dimensions, we avoid mixing representations from various distributions, such as 6D rotational and categorical. The retained features are all 3D continuous representations, ensuring a uniform distribution that aligns better with the diffusion-based generation framework.
To further optimize the motion representations, we then project those essential features into a compact and fine-grained latent space using a motion AutoEncoder (AE). Compared with the motion Variational AutoEncoder (VAE) [12], the deterministic AE projection avoids the variation in the motion latents (i.e. $\boldsymbol{\epsilon}$ incorporation), providing more stable representations that are better suited for diffusion modeling and motion reconstruction. (In Appendix, we show training baseline methods using only essential dimensions can already lead to significant improvements, and processing into latent space may further enhance the results.)
@@
Following previous works [26, 111, 12], the encoder $E$ downsamples $\mathbf{X}$ from $N$ length to $\mathbf{x}$ of $n$ length. The decoder $D$ upsamples it back to $N$ length. With this integration, our method can also use motion latent $\mathbf{x}$ in the diffusion process, which offers acceleration for both training and sampling. In addition, since the downsampling introduces temporal awareness [24], it can improve the temporal coherence of autoregressive diffusion in Sec. 3.2.
@@
where $t$ is continuous timestep and define velocity as:
@@
where $\alpha_{t}$, $\sigma_{t}$ are continous time coefficients, $\dot{\alpha}_{t}=\frac{\mathbf{d}\alpha_{t}}{\mathbf{d}t}$, $\dot{\sigma}_{t}=\frac{\mathbf{d}\sigma_{t}}{\mathbf{d}t}$. We can also deduce score $\mathbf{s}(\mathbf{x},t)$ from $\mathbf{v}(\mathbf{x},t)$.
@@ 3.2 Diffusion: An Autoregressive Approach
Autoregressive methods [26, 75, 76, 111] demonstrate significant advantages in motion generation. Instead of generating the entire sequence $\mathbf{x}^{1:n}$ with a condition c by modeling $p(\mathbf{x}^{1:n}|c)$, autoregressive generation can simplify training and generation into a chain probability:
@@
where each conditional probability $p(\mathbf{x}^{i}|\mathbf{x}^{<i})$ represents the likelihood of generating motion $\mathbf{x}_{i}$ given previous ones $\mathbf{x}^{<i}$, which makes training objective significantly easier.
Naively training diffusion models to perform autoregression using MSE loss fails as it’s simplified to be a regression problem rather than explicitly capturing chained probabilistic distributions of $p(\mathbf{x}^{1:n}|c)$ in autoregressive generation.
Recent advances [51, 91] in image generation demonstrate the potential of autoregressive continuous image modeling. They leverage logits from an autoregressive model as conditioning parameters into a continuous sampling network to better model underlying probability. Inspired by this, we revisit human motion diffusion models from an autoregressive perspective.
@@ 3.2.1 Masked Autoregressive Motion Generation
We follow the masked autoregressive approach proposed by MAR [51]. In each autoregressive iteration, we define unmasked motion latents as $um=\mathbf{x}^{i_{1}:i_{k}}$ and masked motion latents as $m=\mathbf{x}^{j_{1}:j_{n-k}}$, where $k$ is the number of autoregressive steps. The unmasked latents can be refined in a new pseudo-order (a flexible reordering process that can be sequential, random, or any custom ordering) $um=\mathbf{x}^{\prime 1:k}$ to serve as previously generated blocks. The masked motion latents $m=\mathbf{x}^{\prime k+1:n}$ represent the motion latents that need to be generated based on $um$ with condition $c$. Formally, this process is
$p(\mathbf{x}^{\prime 1:n}|c)=p(\text{m}|c)\prod_{j=1}^{k}p(\text{m}|\text{um})\vspace{-1.5mm}$ (13)
We visually illustrate motion masked autoregressive in the right-most part of Fig. 4. Note that because of masked autoregression, our approach is capable of predicting multiple latents, not limited to two shown in Fig. 4.
@@ 3.2.2 Autoregressive Generation Branch
Masked Autoregressive Transformer is designed to process and understand time-variant motion data, and provide rich contextual condition $z$ for the diffusion branch. Given unmasked motion latent sequence $um$ as previous generated motion latents, the masked autoregressive transformer $g$ will produce conditions $z$ to the diffusion branch to generate latents at position of masked motion latents $m$ by:
@@
Given the relative simplicity of motion data compared to image data, we only use one single AdaLN [28] transformer [93] layer. To balance the generation performance and speed, we use bidirectional attention, same as many previous methods [76, 75, 26].
Diffusion MLPs adopts MLPs as its primary structure because with autoregressive modeling, the motion data input into the diffusion branch is independent single $D$-dimensional motion latent. The single $D$-dimensional structure input data aligns well with the MLP’s simplicity and strength in channel-wise manipulation. Unlike MAR [51], which scales the understanding models, we scale generative diffusion MLPs to various model sizes. Using the condition $z$ from Masked Autoregressive Transformer, diffusion branch produce each motion latents $\mathbf{x}^{\prime i}$ in masked token $m$’s position at each timestep $t$ by:
@@
During training, we randomly mask a subset of $k$ motion latents with a learnable continuous mask vector following the cosine masking schedule from MoMask [26]. The autoregressive model learns to provide accurate signals based on the unmasked latents given time step and text condition. The diffusion MLPs utilize this signal to predict $\boldsymbol{\epsilon}$ or $\mathbf{v}(\mathbf{x},t)$ on masked latents. Training objectives for the entire generation branch are denoted as:
@@
For velocity prediction, $\mathcal{L}_{GB}=$
@@
During sampling, given previous latents $um$, we simply add mask vectors to the sequence, allowing the autoregressive model to generate appropriate signals for masked positions. The sampling process can be denoted as:
@@
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ for noise prediction. The velocity prediction uses ODE sampling with step size $\Delta t$:
@@
To address biases in the current evaluation approach shown in Sec. 2.3 due to the unnecessary focus on redundant motion representations, we propose a new evaluation framework that focuses exclusively on essential features (the only ones meaningful), enabling a fairer generation evaluation.
We first construct an evaluator that retains the architecture of the widely used evaluator [24], which consists of a convolutional movement encoder, a GRU [13]-based motion encoder, and a GRU-based text encoder using GloVe [68] embeddings. This evaluator is trained solely on essential dimensions that meaningfully contribute to final motion generation with no other alterations.
@@
By training exclusively on essential dimensions, we ensure the evaluators capture only meaningful features from the final generated motion. The adoption of dual evaluators also provides a more robust and comprehensive framework to accurately compare different-based generation methods. More importantly, all baselines can generate outputs containing essential dimensions with no additional operations.
@@ 4 Experiment
Evaluation Metrics. Following Section 3.3, we employ two evaluators trained with only essential dimensions: one architecturally identical to the one proposed in T2M [24] and a CLIP-based evaluator. Using the T2M evaluator, we adopt evaluation metrics from T2M, including (1) R-Precision (Top-1, Top-2, and Top-3 accuracies) and Matching, which measures the semantic alignment between generated motion embeddings and their corresponding captions’ glove embedding; (2) Fréchet Inception Distance (FID), which assesses the statistical similarity between ground truth and generated motion distributions; and (3) MultiModality, which measures the diversity of generated motion embeddings per same text prompt. Using the CLIP-based evaluator, we include the CLIP-score [30], which measures the compatibility of motion-caption pairs by calculating the cosine similarity between generated motion and its caption.
@@ 4.2 Results and Analysis
Following previous works [24, 88], we conduct each experiment 20 times on both datasets and report the mean result along with a 95% confidence interval. To ensure a fair comparison on the new evaluators, we train all baseline models from scratch following their original methods on the same obtained dataset (full dimension). We present the quantitative results of our method alongside baseline state-of-the-art human motion generation methods in Tab. 5, and the qualitative comparison results in Fig. 5. In addition we also present model scalability results in App. D
As observed, our method achieves superior performance across multiple metrics, including FID, R-Precision, Matching score, and CLIP-score, consistently outperforming baseline methods with non-marginal improvements on both KIT-ML and HumanML3D datasets. Compared to diffusion-based baseline methods, our approach showcases a significantly stronger ability to generate stable motion that follows closely to text instructions and aligns with ground truth. Notably, while the SOTA diffusion-based baseline method ReMoDiffuse [114] relies on additional data retrieval from a large database to achieve high-quality motion generation, our method surpasses ReMoDiffuse’s performance results without requiring auxiliary formulations. In comparison to VQ-based baseline methods, our approach maintains better motion quality and also delivers greater diversity, a quality where VQ methods underperform.
@@ 4.3 Ablation Study
In the ablation study, we further study the impact of reforming data representation and distribution, and autoregressive modeling in our method. We present ablation results in Tab. 6 and an optimization routine in Appendix. The results shows that both proposed components contribute greatly.
@@ 5 Related Work
In this section, we provide a brief overview of related works due to space constraints. We also provide a more detailed version of related works in the Appendix.
VQ-Based Human Motion Generation TM2T [25] first introduces Vector Quantization (VQ) to text-to-human motion generation, enabling discrete motion token modeling. T2M-GPT [111] extended this by leveraging a GPT[6] to motion autoregressive generation. Subsequent methods have sought to integrate a larger model[37, 115] (e.g. large language models), or manipulate attention mechanisms  [119]. Most recently, MMM [76] and MoMask [26] revisit generation methodology by employing bidirectional masked generation techniques inspired by MaskGIT [8]. In this paper, we examine the strengths of these methods and improve a diffusion model with these insights.
Diffusion-Based Human Motion Generation. Inspired by the success of denoising diffusion models in image generation domain [32, 86], several pioneering works [88, 44, 113] have adapted denoising diffusion processes to human motion generation. Building on these works, MLD [12]further optimized the denoising process in latent space to improve training and sampling efficiency. Recent methods have diversified their focus, exploring retrieval-augmention [114], controllable generation [15], as well as investigating advanced architectures [117] such as Mamba [21]. In this paper, we thoroughly investigate the limitations of diffusion-based methods and propose to address them.
Autoregressive Generation with Continous Data. GIVT first introduced the idea of leveraging outputs from an autoregressive model as parameters for a Gaussian Mixture Model, enabling probabilistic autoregressive modeling and generation. MAR then utilized logits from a masked autoregressive model as input to a small diffusion branch, producing more fine-grained generation. Inspired by these approaches, in this paper, we propose a novel framework that integrates diffusion-based motion generation with autoregression to achieve enhanced generative performance.
@@ 6 Conclusion
In conclusion, we introduce a novel diffusion-based generative framework for text-driven 3D human motion generation. Our method reforms motion data representation and distribution to better fit the diffusion model, incorporates masked autoregressive training and sampling techniques and is evaluated by more robust evaluators. Extensive experiments demonstrate our method’s superior generation performance in KIT-ML and HumanML3D datasets.
Acknowledgement.  Yiming Xie was supported by the Apple Scholars in AI/ML PhD fellowship.
@@ Appendix A Detailed Deduction
Appendix B Detailed Related Works
Human Motion Generation. Early text-to-motion approaches [2, 24, 69, 70, 87, 106] attempt to align the latent spaces of text and motion. However, this strategy encounters significant challenges in generating high-fidelity motions due to the inherent difficulty of seamlessly aligning these fundamentally distinct latent spaces. Consequently, recent advancements in human motion generation have shifted focus toward diffusion-based and VQ-based methods, as discussed below.
Diffusion-Based Human Motion Generation. Inspired by the success of denoising diffusion models in the image generation domain [32, 86], several pioneering works [88, 44, 113] have adapted denoising diffusion processes to human motion generation. Building on these works, MLD [12] further optimized the denoising process in latent space to improve training and sampling efficiency. PhysDiff [110] added the physical constraints in the motion generation. And a lot of following works [5, 54, 73, 40, 112, 4, 49, 18, 33, 95, 121, 60] keep exploring diffusion-based human motion generation from different perspectives. In this paper, we thoroughly investigate the limitations of diffusion-based methods and propose a novel approach to address them.
VQ-Based Human Motion Generation. TM2T [25] first introduces Vector Quantization (VQ) to text-to-human motion generation, enabling discrete motion token modeling. A lot of the following works [111, 109, 76, 26, 8, 75, 119, 53] improved the VQ-based methods. T2M-GPT [111] extended this by leveraging a GPT [6] to motion autoregressive generation. Subsequent methods have sought to integrate a larger model [37, 115] (e.g. large language models), or manipulate attention mechanisms  [119]. Most recently, MMM [76] and MoMask [26] revisit generation methodology by employing bidirectional attention-based masked generation techniques inspired by MaskGIT [8]. BAMM [75] introduced a dual-iteration framework that combines unidirectional generation with bidirectional refinement to enhance the coherence of generated motions. The concurrent work ScaMo [61] explored the scaling law in human motion generation by training the model with large-scale data. In this paper, we examine the strengths of these approaches and improve a diffusion model inspired by these insights.
Autoregressive Generation with Continous Data. In motion synthesis, recent works [116, 11, 85, 22, 89] have started to explore integrating autoregressive structures into diffusion-based frameworks. However, due to the challenges of performing direct causal next motion prediction with MSE loss (as done in discrete token settings), these methods typically only use previously generated motion as a prefix condition, rather than modeling the next step motion directly using previous motion as input. In contrast, recent image generation methods have explored tighter coupling between autoregression and diffusion. GIVT introduced the idea of giving previous generation as input, using outputs from an autoregressive model as parameters for a Gaussian Mixture Model to enable probabilistic chaining of autoregressive generation. MAR further refined this by feeding logits from a masked autoregressive model into a small diffusion branch, producing more fine-grained generation. Inspired by these approaches, we propose to integrate diffusion-based motion generation with masked autoregression, enabling a more direct autoregressive technique beyond simple prefix conditioning to achieve improved generative performance.
Human Motion Generation and Beyond. Recent methods have diversified their focus, exploring retrieval-augmention [114], controllable generation [101, 15, 43, 41, 79, 94, 74], human-scene/object interactions [67, 34, 46, 104, 72, 48, 10, 97, 20, 108, 105, 63, 14, 100, 38, 50, 59, 17, 118, 96], human-human interaction [36, 103, 99, 19, 56, 7], stylized human motion generation [120, 27, 52], more datasets [102, 58], long-motion generation [122, 71], voice-conditioned motion generation [9], unified motion generation and understanding [47], shape-aware motion generation [90], fine-grained text controlled generation [123, 35, 107, 84, 39], fine-tuning pretrained motion generation model as priors [42, 83], as well as investigating advanced architectures [117, 98] such as Mamba [21].